# Improving LLM Reasoning with Multi-Agent Tree-of-Thought Validator Agent

### v1

@@ Abstract
Multi-agent strategies have emerged as a promising approach to enhance the reasoning abilities of Large Language Models (LLMs) by assigning specialized roles in the problem-solving process. Concurrently, Tree of Thoughts (ToT) methods have shown potential in improving reasoning for complex question-answering tasks by exploring diverse reasoning paths. A critical limitation in multi-agent reasoning is the ’Reasoner’ agent’s shallow exploration of reasoning paths. While ToT strategies could help mitigate this problem, they may generate flawed reasoning branches, which could harm the trustworthiness of the final answer. To leverage the strengths of both multi-agent reasoning and ToT strategies, we introduce a novel approach combining ToT-based Reasoner agents with a Thought Validator agent. Multiple Reasoner agents operate in parallel, employing ToT to explore diverse reasoning paths. The Thought Validator then scrutinizes these paths, considering a Reasoner’s conclusion only if its reasoning is valid. This method enables a more robust voting strategy by discarding faulty reasoning paths, enhancing the system’s ability to tackle tasks requiring systematic and trustworthy reasoning. Our method demonstrates superior performance compared to existing techniques when evaluated on the GSM8K dataset, outperforming the standard ToT strategy by an average 5.6% across four LLMs.
@@ 1 Introduction
LLMs have demonstrated remarkable capabilities across various tasks, yet they often struggle with complex reasoning, particularly in situations where human-like reasoning capabilities are crucial Zelikman et al. (2023). To address this, multi-agent strategies have emerged as a promising method to enhance LLM reasoning. Using multi-agent strategies, multiple specialized agents collaborate, with each agent assigned distinct roles in the problem-solving process. By allowing different agents to tackle various aspects of a task, we they are able to utilize specialized expertise to each phase of the task to improve performance Guo et al. (2024). This has been shown to improve the quality of answers in reasoning-intensive domains. However, despite the promise of multi-agent reasoning, one critical limitation remains: Reasoner agents often explore reasoning paths shallowly, failing to fully consider the complexity of the problem space. Tree of Thoughts (ToT) methods offer a potential solution to this limitation by encouraging a more systematic exploration of multiple reasoning paths. ToT allows LLMs to simulate human-like thought processes by branching out and examining various possibilities before converging on a solution Yao et al. (2023). By enabling LLMs to consider diverse reasoning pathways, ToT can mitigate the shallow exploration issue seen in some other multi-agent systems. However, while ToT encourages exploration, it also introduces a new challenge: the risk of generating flawed reasoning branches. Without proper validation, these erroneous paths could lower the overall trustworthiness of the final answer.
To address these challenges, we propose a novel approach that combines the strengths of multi-agent reasoning with ToT while introducing a critical validation mechanism. In our framework, multiple Reasoner agents operate in parallel, each employing ToT to explore different reasoning paths. These Reasoner agents are supported by a Thought Validator agent, which evaluates the proposed reasoning branches produced by the Reasoners. The Validator discards faulty reasoning branches, ensuring that only logically sound paths contribute to the final decision. This approach allows for both exploration of the problem space and increased reliability of the answers by eliminating flawed reasoning paths before they can impact the outcome. Our proposed approach is evaluated on the GSM8K dataset Cobbe et al. , a benchmark known for its challenging arithmetic reasoning tasks. Results show that our method outperforms existing techniques, demonstrating improved accuracy and trustworthiness in solving complex reasoning problems.
@@ 2.1 Multi-agent Systems for Enhancing LLM Reasoning
By distributing tasks among multiple agents, multi-agent systems aim to improve performance on reasoning-based tasks Du et al. ; Talebirad and Nadiri . For example, CausalGPT Tang et al. introduces evaluative layers to verify the reasoning branches produced by LLMs, while the Counterfactual Multi-Agent Debate (CFMAD) framework Fang et al. provides an innovative method to mitigate the potentially biased reasoning branches of LLMs by assigning agents to fixed roles to generate justifications from specific perspectives, and a third-party judge evaluates these arguments to decide the most rational outcome. Despite these advancements, current methods still suffer from shallow sampling of reasoning paths or majority vote schemes. These techniques can overlook critical inferential errors and are especially prone to early-stage errors, which can propagate through multiple rounds of reasoning. This limitation is especially problematic in complex scenarios where systematically evaluating and eliminating incorrect options is crucial. Recent research has demonstrated that LLMs can effectively identify both factual and inferential mistakes Li et al. , making the integration of a dedicated verification component in multi-agent systems particularly beneficial for assessing the faithfulness and reliability of generated solutions.
@@ 2.2 The Role of the ’Reasoner’ Agent in Multi-Agent Frameworks
Within multi-agent architectures, the Reasoner agent plays a pivotal role. It serves as the system’s core decision-maker, ensuring that valid conclusions are derived from the reasoning process. However, Reasoner agents in current frameworks often struggle to systematically evaluate and eliminate incorrect reasoning paths, particularly in more challenging problem spaces. This bottleneck highlights the need for more advanced reasoning strategies to be integrated into the Reasoner agent. CFMAD has also shown that checking all available options can enhance the overall ability of the multi-agent systems Fang et al. .
@@ 3.1 Reasoner Agent
The Reasoner agents in our multi-agent architecture employ the ToT strategy, which enables structured exploration of reasoning paths in parallel. ToT improves upon Chain of Thought (CoT) prompting Wei et al. by enabling parallel exploration and dynamic path evaluation. While CoT follows a single, linear path, ToT actively explores and evaluates multiple reasoning paths, making it better suited for complex problems that benefit from diverse thought exploration Yao et al. . We formalize the reasoning process as a search over a tree of states. Let $Q$ denote the input prompt or query, and each Reasoner agent $R_{i}$ constructs a Tree of Thoughts $T_{i}(Q)$, where each node represents a state $s_{t}$, which is a distinct point along a reasoning path. A state $s_{t}$ consists of the problem $Q$ and a sequence of intermediate reasoning steps up to that point $z_{1},z_{2},\dots,z_{t}$, with each step $z_{j}$ being a coherent unit of reasoning generated by the language model.
$s_{t}=[Q,z_{1},z_{2},\dots,z_{t}]$
@@ 4 Experiments
Dataset: GSM8K Cobbe et al. is a dataset of 8.5K high-quality linguistically diverse grade school math word problems created by human problem writers. GSM8K is widely recognized as a benchmark for testing arithmetic reasoning in LLMs. The dataset comprises complex, multi-step mathematical word problems that challenge both the reasoning and computation capabilities of LLMs. Our experiments utilized a random subset of 500 samples from the GSM8K dataset as the test set. Following other works on LLM reasoning on the GSM8K dataset, we evaluated the performance of reasoning approaches using accuracy as the primary metric Yao et al. (2023).
@@
Implementation Details: Our experiments cover two versions of OpenAI’s GPT models and two versions of Meta’s Llama 3.1 models Dubey et al. (2024). Specifically, we use GPT-3.5 Turbo OpenAI (2024b) and GPT-4o-mini OpenAI (2024a) from OpenAI, accessed through their API. For the Llama 3.1 models, we employ the 8B NousResearch (2024b) and 70B NousResearch (2024a) parameter variants. These models offer a range of capabilities and sizes, allowing us to explore different trade-offs between model complexity and performance in our experiments. We conduct all of our experiments on four Nvidia DGX A100 80 GB GPUs, and running all these experiments in parallel took about 18 hours. For our baseline comparisons, we employed several prompting strategies. We began with input-output (IO) prompting, a standard approach that transforms a problem input into an output by conditioning on task instructions. We then implemented more advanced techniques, including Chain of Thought (CoT) Wei et al. and the original ToT strategy Yao et al. (2023). For the ToT implementation, we followed the parameters used by  Yao et al. (2023) on the GSM8K dataset, setting a tree depth of 2 and a width of 5. To ensure consistency across our baseline models, we used a temperature of 1 and a top_p value of 1 for IO, CoT, and ToT. However, for our novel Thought Validator Agent, we adjusted these parameters to a temperature of 0.5 and a top_p of 0.4. This adjustment was made to increase the determinism of the Thought Validator’s outputs, as its role is to validate existing reasoning rather than generate creative solutions.
Experimental Results:
@@
Table 1 shows the performance comparison of these different methods. Our results show that our proposed multi-agent ToT reasoner with a Thought Validator agent outperforms the other reasoning methods, showing an improvement of 8.8 percentage points over ToT for GPT-3.5-turbo (from 75.4% to 84.2%). We also see that while ToT and other techniques showed significant improvements over standard IO prompting when the LLM struggled with a task (such as with GPT 3.5 Turbo and Llama 3.1 8B), the performance gap narrowed considerably for problems where the model with standard IO prompting already exhibited strong capabilities (such as with GPT 4o mini, and Llama 3.1 70B). This observation suggests that the efficacy of ToT may be dependent on the complexity of the task and capability of the model, with its benefits more pronounced in challenging reasoning tasks that push the boundaries of the model’s baseline abilities. The effectiveness of ToT in these scenarios can be likened to a teacher providing feedback to a struggling student, guiding them through complex problems, and reinforcing correct thought processes.
@@ 5 Limitations and Conclusion
While the ToT approach has shown promise in enhancing reasoning capabilities, our observations of the outputs and reasoning trees revealed several limitations that warrant further investigation. A key challenge we observed is the lack of dynamic exploration in the search space. The ToT method proposed by Yao et al. (2023) employs a fixed width and depth for the tree structure, which our analysis showed can lead to suboptimal performance in certain scenarios. For instance, when examining the reasoning trees for problems that could be solved efficiently without extensive reasoning, we found that the predetermined depth of exploration often introduced unnecessary complexity, potentially leading to errors or confusion in the reasoning process. Conversely, for problems requiring more in-depth analysis, we observed that the fixed depth proved insufficient, limiting the model’s ability to fully explore complex reasoning paths. Additionally, our proposed approach, while addressing some of these limitations, is computationally expensive due to the use of the ToT method, which requires significant resources for generating and evaluating multiple thought paths.
In conclusion, we have presented a novel approach that combines the ToT strategy with a multi-agent reasoning framework enhanced by a Thought Validator agent. Our method addresses key limitations in existing reasoning strategies for LLMs by enabling a more systematic exploration of reasoning paths while simultaneously improving the reliability of generated solutions. Experimental results on the GSM8K dataset demonstrate that our approach outperforms state-of-the-art methods, particularly for complex arithmetic reasoning tasks.
@@ 6 Social Impact Statement
Reasoner Reasoning Summary Final Answer Verified
R1 Incorrect reasoning. Algebraic error leads to the incorrect conclusion that Tasha did not mow Joe’s lawn. \$80 False
R2 Correct strategy, but incorrect total earnings from mowing lawns (\$60x instead of \$48x). \$80 False
R3 Accurate reasoning. Correctly calculates the total earnings from mowing lawns and finds lemonade income to be \$26. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified
R1 Repeated the previous error, miscalculating Tasha’s income from lawn mowing. \$80 False
R2 Corrected earlier miscalculation but again used the wrong mowing total, leading to an incorrect conclusion. \$80 False
R3 Maintained the correct reasoning and final answer (\$26) as in Round 1. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified
R1 Corrects earlier algebraic error, but still does not address the lemonade sales correctly. \$80 False
R2 Adjusts previous mistake but there is a slight issue in the final calculation but the Validator Agent was not able to detect it. \$32 False
R3 Maintained the correct reasoning and final answer (\$26) as in Round 1. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified
R1 Accurate breakdown of laundry items per room and multiplication across 80 rooms. Correct total of 1200 pieces of laundry. 1200 True
R2 CCorrect approach, but overestimated the number of pillowcases, leading to an incorrect total of 1280 pieces of laundry. 1280 False
R3 Correct breakdown of laundry per room, yielding the correct total of 1200 pieces of laundry. 1200 True

### v2

@@ Abstract
Multi-agent strategies have emerged as a promising approach to enhance the reasoning abilities of Large Language Models (LLMs) by assigning specialized roles in the problem-solving process. Concurrently, Tree of Thoughts (ToT) methods have shown potential in improving reasoning for complex question-answering tasks by exploring diverse reasoning paths. A critical limitation in multi-agent reasoning is the ’Reasoner’ agent’s shallow exploration of reasoning paths. While ToT strategies could help mitigate this problem, they may generate flawed reasoning branches, which could harm the trustworthiness of the final answer. To leverage the strengths of both multi-agent reasoning and ToT strategies, we introduce a novel approach combining ToT-based Reasoner agents with a Thought Validator agent. Multiple Reasoner agents operate in parallel, employing ToT to explore diverse reasoning paths. The Thought Validator then scrutinizes these paths, considering a Reasoner’s conclusion only if its reasoning is valid. This method enables a more robust voting strategy by discarding faulty reasoning paths, enhancing the system’s ability to tackle tasks requiring systematic and trustworthy reasoning. Our method demonstrates superior performance compared to existing techniques when evaluated on the GSM8K dataset, outperforming the standard ToT strategy by an average 5.6% across four LLMs. The code and related content can be found in: https://github.com/SecureAIAutonomyLab/MA-ToT
@@ 1 Introduction
LLMs have demonstrated remarkable capabilities across various tasks, yet they often struggle with complex reasoning, particularly in situations where human-like reasoning capabilities are crucial [18]. To address this, multi-agent strategies have emerged as a promising method to enhance LLM reasoning. Using multi-agent strategies, multiple specialized agents collaborate, with each agent assigned distinct roles in the problem-solving process. By allowing different agents to tackle various aspects of a task, we they are able to utilize specialized expertise to each phase of the task to improve performance [5]. This has been shown to improve the quality of answers in reasoning-intensive domains. However, despite the promise of multi-agent reasoning, one critical limitation remains: Reasoner agents often explore reasoning paths shallowly, failing to fully consider the complexity of the problem space. Tree of Thoughts (ToT) methods offer a potential solution to this limitation by encouraging a more systematic exploration of multiple reasoning paths. ToT allows LLMs to simulate human-like thought processes by branching out and examining various possibilities before converging on a solution [17]. By enabling LLMs to consider diverse reasoning pathways, ToT can mitigate the shallow exploration issue seen in some other multi-agent systems. However, while ToT encourages exploration, it also introduces a new challenge: the risk of generating flawed reasoning branches. Without proper validation, these erroneous paths could lower the overall trustworthiness of the final answer.
To address these challenges, we propose a novel approach that combines the strengths of multi-agent reasoning with ToT while introducing a critical validation mechanism. In our framework, multiple Reasoner agents operate in parallel, each employing ToT to explore different reasoning paths. These Reasoner agents are supported by a Thought Validator agent, which evaluates the proposed reasoning branches produced by the Reasoners. The Validator discards faulty reasoning branches, ensuring that only logically sound paths contribute to the final decision. This approach allows for both exploration of the problem space and increased reliability of the answers by eliminating flawed reasoning paths before they can impact the outcome. Our proposed approach is evaluated on the GSM8K dataset [1], a benchmark known for its challenging arithmetic reasoning tasks. Results show that our method outperforms existing techniques, demonstrating improved accuracy and trustworthiness in solving complex reasoning problems.
@@ 2.1 Multi-agent Systems for Enhancing LLM Reasoning
Recent work has demonstrated various approaches to enhance LLM outputs. Over-generation and reranking strategies have shown promising results, with RANKGEN [6] introducing a ranking model that scores generated outputs based on relevance and coherence, and Faithfulness-Aware Decoding [14] demonstrating how different decoding strategies affect output quality. While these methods are effective for general text generation, complex reasoning tasks often benefit from more structured approaches. Multi-agent systems have emerged as a particularly effective framework, improving performance on reasoning-based tasks by distributing tasks among specialized agents [2, 12]. For example, CausalGPT [13] introduces evaluative layers to verify the reasoning branches produced by LLMs, while the Counterfactual Multi-Agent Debate (CFMAD) framework [4] provides an innovative method to mitigate the potentially biased reasoning branches of LLMs by assigning agents to fixed roles to generate justifications from specific perspectives, and a third-party judge evaluates these arguments to decide the most rational outcome. Despite these advancements, current methods still suffer from shallow sampling of reasoning paths or majority vote schemes. These techniques can overlook critical inferential errors and are especially prone to early-stage errors, which can propagate through multiple rounds of reasoning. This limitation is especially problematic in complex scenarios where systematically evaluating and eliminating incorrect options is crucial. Recent research has demonstrated that LLMs can effectively identify both factual and inferential mistakes [7], making the integration of a dedicated verification component in multi-agent systems particularly beneficial for assessing the faithfulness and reliability of generated solutions.
@@ 2.2 The Role of the ’Reasoner’ Agent in Multi-Agent Frameworks
Within multi-agent architectures, the Reasoner agent plays a pivotal role. It serves as the system’s core decision-maker, ensuring that valid conclusions are derived from the reasoning process. However, Reasoner agents in current frameworks often struggle to systematically evaluate and eliminate incorrect reasoning paths, particularly in more challenging problem spaces. This bottleneck highlights the need for more advanced reasoning strategies to be integrated into the Reasoner agent. CFMAD has also shown that checking all available options can enhance the overall ability of the multi-agent systems [4].
@@ 3.1 Reasoner Agent
The Reasoner agents in our multi-agent architecture employ the ToT strategy, which enables structured exploration of reasoning paths in parallel. ToT improves upon Chain of Thought (CoT) prompting [15] by enabling parallel exploration and dynamic path evaluation. While CoT follows a single, linear path, ToT actively explores and evaluates multiple reasoning paths, making it better suited for complex problems that benefit from diverse thought exploration [16]. We formalize the reasoning process as a search over a tree of states. Let $Q$ denote the input prompt or query, and each Reasoner agent $R_{i}$ constructs a Tree of Thoughts $T_{i}(Q)$, where each node represents a state $s_{t}$. A state $s_{t}=[Q,z_{1},z_{2},\dots,z_{t}]$ consists of the problem $Q$ and a sequence of intermediate reasoning steps up to that point $z_{1},z_{2},\dots,z_{t}$, with each step $z_{j}$ being a coherent unit of reasoning generated by the language model.
@@ 4 Experiments
Dataset: GSM8K [1] is a dataset of 8.5K high-quality linguistically diverse grade school math word problems created by human problem writers. GSM8K is widely recognized as a benchmark for testing arithmetic reasoning in LLMs. The dataset comprises complex, multi-step mathematical word problems that challenge both the reasoning and computation capabilities of LLMs. Our experiments utilized a random subset of 500 samples from the GSM8K dataset as the test set. Following other works on LLM reasoning on the GSM8K dataset, we evaluated the performance of reasoning approaches using accuracy as the primary metric [17].
@@
Implementation Details: Our experiments cover two versions of OpenAI’s GPT models and two versions of Meta’s Llama 3.1 models [3]. Specifically, we use GPT-3.5-turbo-0125 [11] and GPT-4o-mini-2024-07-18 [10] from OpenAI, accessed through their API. For the Llama 3.1 models, we employ the 8B [9] and 70B [8] parameter variants. These models offer a range of capabilities and sizes, allowing us to explore different trade-offs between model complexity and performance in our experiments. We conduct all of our experiments on four Nvidia DGX A100 80 GB GPUs, and running all these experiments in parallel took about 18 hours. For our baseline comparisons, we employed several prompting strategies. We began with input-output (IO) prompting, a standard approach that transforms a problem input into an output by conditioning on task instructions. We then implemented more advanced techniques, including Chain of Thought (CoT) [15] and the original ToT strategy [17]. For the ToT implementation, we followed the parameters used by  Yao et al. [17] on the GSM8K dataset, setting a tree depth of 2 and a width of 5. To ensure consistency across our baseline models, we used a temperature of 1 and a top_p value of 1 for IO, CoT, and ToT. However, for our novel Thought Validator Agent, we adjusted these parameters to a temperature of 0.5 and a top_p of 0.4. This adjustment was made to increase the determinism of the Thought Validator’s outputs, as its role is to validate existing reasoning rather than generate creative solutions.
@@
Experimental Results: Table 1 shows the performance comparison of these different methods. Our results show that our proposed multi-agent ToT reasoner with a Thought Validator agent outperforms the other reasoning methods, showing an improvement of 8.8 percentage points over ToT for GPT-3.5-turbo (from 75.4% to 84.2%). We also see that while ToT and other techniques showed significant improvements over standard IO prompting when the LLM struggled with a task (such as with GPT 3.5 Turbo and Llama 3.1 8B), the performance gap narrowed considerably for problems where the model with standard IO prompting already exhibited strong capabilities (such as with GPT 4o mini, and Llama 3.1 70B). This observation suggests that the efficacy of ToT may be dependent on the complexity of the task and capability of the model, with its benefits more pronounced in challenging reasoning tasks that push the boundaries of the model’s baseline abilities. The effectiveness of ToT in these scenarios can be likened to a teacher providing feedback to a struggling student, guiding them through complex problems, and reinforcing correct thought processes.
@@ 5 Limitations and Conclusion
While the ToT approach has shown promise in enhancing reasoning capabilities, our observations of the outputs and reasoning trees revealed several limitations that warrant further investigation. A key challenge we observed is the lack of dynamic exploration in the search space. The ToT method proposed by Yao et al. [17] employs a fixed width and depth for the tree structure, which our analysis showed can lead to suboptimal performance in certain scenarios. For instance, when examining the reasoning trees for problems that could be solved efficiently without extensive reasoning, we found that the predetermined depth of exploration often introduced unnecessary complexity, potentially leading to errors or confusion in the reasoning process. Conversely, for problems requiring more in-depth analysis, we observed that the fixed depth proved insufficient, limiting the model’s ability to fully explore complex reasoning paths. Additionally, our proposed approach, while addressing some of these limitations, is computationally expensive due to the use of the ToT method, which requires significant resources for generating and evaluating multiple thought paths. Our analysis shows that our multi-agent ToT approach with the Thought Validator requires substantially more compute resources than standard methods. In our 500-sample evaluation, the average token usage per question increases significantly: for GPT-3.5-turbo, from 256 tokens with CoT to 4000 tokens with ToT, while for GPT-4o-mini, from 341 to 10,600 tokens. Each Reasoner agent requires approximately 20 API calls per problem, with our multi-agent approach multiplying this cost by the number of agents plus validation steps. While this increased computational investment yields meaningful improvements in reasoning accuracy—demonstrated by an 8.8 percentage point improvement for GPT-3.5-turbo—the trade-off may require careful consideration in resource-constrained environments. Furthermore, while our evaluation on GSM8K demonstrates the effectiveness of our approach for arithmetic reasoning, testing on additional reasoning-intensive benchmarks would help establish the method’s generalizability across different types of reasoning tasks.
The Thought Validator agent demonstrates strong capabilities in assessing reasoning paths, and its effectiveness could be further enhanced through advanced validation techniques such as ensemble validation strategies or meta-learning approaches to improve robustness across diverse reasoning scenarios. Furthermore, while our evaluation on GSM8K demonstrates the effectiveness of our approach for mathematical reasoning, testing on additional reasoning-intensive benchmarks would help establish the method’s generalizability across different types of reasoning tasks.
In conclusion, we have presented a novel approach that combines the ToT strategy with a multi-agent reasoning framework enhanced by a Thought Validator agent. Our method addresses key limitations in existing reasoning strategies for LLMs by enabling a more systematic exploration of reasoning paths while simultaneously improving the reliability of generated solutions. Experimental results on the GSM8K dataset demonstrate that our approach outperforms state-of-the-art methods, particularly for complex arithmetic reasoning tasks. Future work could explore dynamic tree structuring based on problem complexity, potentially improving efficiency and performance across a wider range of problem types.
@@ 6 Social Impact Statement
Reasoner Reasoning Summary Final Answer Verified R1 Incorrect reasoning. Algebraic error leads to the incorrect conclusion that Tasha did not mow Joe’s lawn. \$80 False R2 Correct strategy, but incorrect total earnings from mowing lawns (\$60x instead of \$48x). \$80 False R3 Accurate reasoning. Correctly calculates the total earnings from mowing lawns and finds lemonade income to be \$26. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified R1 Repeated the previous error, miscalculating Tasha’s income from lawn mowing. \$80 False R2 Corrected earlier miscalculation but again used the wrong mowing total, leading to an incorrect conclusion. \$80 False R3 Maintained the correct reasoning and final answer (\$26) as in Round 1. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified R1 Corrects earlier algebraic error, but still does not address the lemonade sales correctly. \$80 False R2 Adjusts previous mistake but there is a slight issue in the final calculation but the Validator Agent was not able to detect it. \$32 False R3 Maintained the correct reasoning and final answer (\$26) as in Round 1. \$26 True
@@
Reasoner Reasoning Summary Final Answer Verified R1 Accurate breakdown of laundry items per room and multiplication across 80 rooms. Correct total of 1200 pieces of laundry. 1200 True R2 CCorrect approach, but overestimated the number of pillowcases, leading to an incorrect total of 1280 pieces of laundry. 1280 False R3 Correct breakdown of laundry per room, yielding the correct total of 1200 pieces of laundry. 1200 True